OpenStudy (anonymous):

Which value for x makes the sentence true? 16.03 + x = 63.83 47.69 47.8 79.86 80.86

OpenStudy (anonymous):

help please

OpenStudy (alexandervonhumboldt2):

subtract 16.03 from both sides: 16.03+x-16.03=63.83-16.03 x=47.8

OpenStudy (anonymous):

solve or trial and error...

OpenStudy (anonymous):

thanks i got afey more

OpenStudy (babylove18):

63.83-16.03=47.8

OpenStudy (anonymous):

Which value for x makes the sentence true? 5,810 + x = 6,779

OpenStudy (babylove18):

6779-5810=969

OpenStudy (anonymous):

Which value for x makes the sentence true? 33.88 + x = 39.46 5.68 72.24 73.34 5.58

OpenStudy (anonymous):

need help

OpenStudy (babylove18):

5.58 @mjmahmood

OpenStudy (anonymous):

thanks

OpenStudy (babylove18):

your welcome